# DACAD: Domain Adaptation Contrastive Learning for Anomaly Detection in Multivariate Time Series

### v1

@@ 4.3 Baselines Comparison
To assess our model against ContextDA Lai et al. (2023) -the only exiting UDA for TAD model-, we use the same datasets and metrics reported in ContextDA’s main paper, as its code is unavailable. On the SMD dataset, ContextDA achieves an average macro F1 of 0.63 and AUROC of 0.75, whereas our model achieves a higher average macro F1 of 0.81 and AUROC of 0.86. Similarly, on the Boiler dataset, ContextDA’s average macro F1 is 0.50 and AUROC 0.65, compared to our model’s superior performance with an average macro F1 of 0.63 and AUROC of 0.71. Details of these results are provided in the appendix, due to page limitations.
4.4 Ablation Study
@@
Table 3: Effect of loss components on MSL dataset (source: F-5).
@@
Table 3 offers several key insights into the impact of different loss components on the MSL dataset using F-5 as a source. Removing the target self-supervised CL (w/o $\mathcal{L}_{\text{SelfCont}}$) leads to lower metrics (F1: 0.481, AUPR: 0.495, AUROC: 0.697). Moreover, excluding source supervised CL (w/o $\mathcal{L}_{\text{SupCont}}$) reduces effectiveness (F1: 0.463, AUPR: 0.427, AUROC: 0.639), highlighting its role in capturing source-specific features which are crucial for the model’s overall accuracy. Similarly, omitting the discriminator component results in performance reduction (F1: 0.471, AUPR: 0.484, AUROC: 0.699). However, the most significant decline occurs without the classifier (F1: 0.299, AUPR: 0.170, AUROC: 0.503), underscoring its crucial role in effectively distinguishing between normal/anomaly classes. Overall, the best results (F1: 0.595, AUPR: 0.554, AUROC: 0.787) are achieved with all components, highlighting the effectiveness of an integrated approach.
@@
To elucidate the effectiveness of UDA within DACAD, we examine the feature representations from the MSL dataset, as illustrated in Figure 2. It presents the t-SNE 2D embeddings of DACAD feature representations $\phi^{CL}(\phi^{R}(w))$ for MSL dataset. Each point represents a time series window, which can be normal, anomalous or anomaly-injected. These representations highlight the domain discrepancies between source and target entities and demonstrate how DACAD aligns the time series window features effectively. The comparison of feature representations with and without UDA reveals a significant domain shift when UDA is not employed, between source (entity F-5) and target (entity T-5) within MSL dataset.
@@
Effect of CEC: Table 4 compares the performance of our CEC classifier with two BCE and DeepSVDD - on the MSL dataset using F-5 as a source. Our proposed CEC shows superior performance compared to BCE and DeepSVDD across three metrics on the MSL dataset. With the highest F1 score, it demonstrates a better balance of precision and recall. Its leading performance in AUPR indicates greater effectiveness in identifying positive classes in imbalanced datasets. Additionally, CEC’s higher AUROC suggests it is more capable of distinguishing between classes.
Table 4: Effect of CEC classifier on MSL dataset (source: F-5).
@@
Table 5: Effect of anomaly injection on MSL dataset (source: F-5).
@@
Effect of anomaly injection: We study the impact of anomaly injection in Table 5, on MSL dataset when using F-5 as a source. It shows that anomaly injection significantly improves all metrics (F1: 0.595, AUPR: 0.554, AUROC: 0.787), enhancing the model’s ability to differentiate between normal and anomalous patterns, thereby improving DACAD’s overall accuracy. Without anomaly injection, there’s a notable decline in performance, emphasizing its role in precision. The higher standard deviation in AUROC scores without injection suggests more variability and less stability in the model’s performance. The study underscores the vital role of anomaly injection in improving anomaly detection models. It reveals that incorporating anomaly injection not only boosts detection accuracy but also enhances the model’s overall stability and reliability.
5 Conclusion
@@
The DACAD model stands as a notable innovation in the field of TAD, particularly effective in environments with limited labelled data. By melding domain adaptation and contrastive learning, it applies labelled data from one domain to detect anomalies in another. Its anomaly injection mechanism, introducing a spectrum of synthetic anomalies, significantly bolsters the model’s adaptability and robustness across various domains. Our evaluations on diverse real-world datasets establish DACAD’s superiority in handling domain shifts and outperforming existing models. Its capability to generalise and accurately detect anomalies, regardless of the scarcity of labelled data in the target domain, represents a significant contribution to TAD. In future work, we aim to refine the anomaly injection process further, enhancing the model’s ability to simulate a broader range of anomalous patterns. Additionally, we plan to evolve the model to encompass univariate time series analysis, broadening its scope and utility.
Appendix A Datasets’ Descriptions
Mars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) [^2] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl datasets Hundman et al. (2018) are real-world datasets collected from NASA spacecraft. These datasets contain anomaly information derived from reports of incident anomalies for a spacecraft monitoring system. MSL and SMAP comprises of 27 and 55 datasets respectively and each equipped with a predefined train/test split, where, unlike other datasets, their training set are unlabeled.
Server Machine Dataset (SMD) [^3] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset Su et al. (2019) is gathered from 28 servers, incorporating 38 sensors, over a span of 10 days. During this period, normal data was observed within the initial 5 days, while anomalies were sporadically injected during the subsequent 5 days. The dataset is also equipped with a predefined train/test split, where the training data is unlabeled.
Boiler Fault Detection Dataset. [^4] https://github.com/DMIRLAB-Group/SASA-pytorch/tree/main/datasets/Boiler Cai et al. (2021) The Boiler dataset includes sensor information from three separate boilers, with each boiler representing an individual domain. The objective of the learning process is to identify the malfunctioning blowdown valve in each boiler. Obtaining samples of faults is challenging due to their scarcity in the mechanical system. Therefore, it’s crucial to make use of both labelled source data and unlabeled target data to enhance the model’s ability to generalise.
Appendix B Baselines
Below, we will provide an enhanced description of the UDA models for time series classification and anomaly detection.
Additionally, we provide a description of the four prominent and state-of-the-art TAD models that were used for comparison with DACAD. We have selected the models from different categories of TAD, namely, unsupervised reconstruction-based (OmniAnomaly Su et al. (2019) model, unsupervised forecasting-based (THOC Shen et al. (2020)), and self-supervised contrastive learning (TS2Vec Yue et al. (2022) and DCdetector Yang et al. (2023)) TAD models.
VRADA [^5] https://github.com/floft/vrada combines deep domain confusion tzeng2014deep with variational recurrent adversarial deep domain adaptation Purushotham et al. (2016), which simultaneously optimises source domain label prediction, MMD and domain discrimination with the latent representations generated by the LSTM encoder. Meanwhile, the reconstruction objective of AE-LSTM is performed to detect anomalies.
CLUDA [^6] https://github.com/oezyurty/CLUDA is a novel framework developed for UDA in time series data and evaluated on time series classification. This framework utilises a contrastive learning approach to learn domain-invariant semantics within multivariate time series data, aiming to preserve label information relevant to prediction tasks. CLUDA is distinctive as it is the first UDA framework designed to learn contextual representations of time series data, maintaining the integrity of label information. Its effectiveness is demonstrated through evaluations of various time series classification datasets.
ContextDA [^7] There is no implementation available for this model, so we rely on the results claimed in the paper. is a sophisticated approach for detecting anomalies in time series data. It employs context sampling formulated as a Markov decision process and uses deep reinforcement learning to optimise the domain adaptation process. This model is designed to generate domain-invariant features for better anomaly detection across various domains. It has shown promise in transferring knowledge between similar or entirely different domains.
OmniAnomaly [^8] https://github.com/smallcowbaby/OmniAnomaly is a model operating on an unsupervised basis, employing a Variational Autoencoder (VAE) to handle multivariate time series data. It identifies anomalies by evaluating the reconstruction likelihood of specific data windows.
THOC [^9] We utilised the authors’ shared implementation, as it is not publicly available. utilises a multi-layer dilated recurrent neural network (RNN) alongside skip connections in order to handle contextual information effectively. It adopts a temporal hierarchical one-class network approach for detecting anomalies.
TS2Vec [^10] https://github.com/yuezhihan/ts2vec is an unsupervised model that is capable of learning multiple contextual representations of MTS and UTS semantically at various levels. This model employs contrastive learning in a hierarchical way, which provides a contextual representation. A method within TS2Vec has been proposed for application in TAD.
DCdetector. [^11] https://github.com/DAMO-DI-ML/KDD2023-DCdetector is distinctive for its use of a dual attention asymmetric design combined with contrastive learning. Unlike traditional models, DCdetector does not rely on reconstruction loss for training. Instead, it utilises pure contrastive loss to guide the learning process. This approach enables the model to learn a permutation invariant representation of time series anomalies, offering superior discrimination abilities compared to other methods.
Table 6: Macro F1/AUROC results on SMD and Boiler dataset. The most optimal evaluation results are displayed in bold, while the second best ones are indicated by underline.
Data Models (Macro F1/AUROC)
src $\mapsto$ trg AE-MLP AE_LSTM RDC VRADA SASA ContexTDA DACAD
SMD 1-1 $\mapsto$ 1-2 0.72 / 0.83 0.74 / 0.90 0.74 / 0.89 0.74 / 0.91 0.59 / 0.63 0.75 / 0.91 0.80 / 0.84
1-1 $\mapsto$ 1-3 0.57 / 0.70 0.49 / 0.41 0.57 / 0.72 0.49 / 0.41 0.61 / 0.90 0.57 / 0.75 0.71 / 0.84
1-1 $\mapsto$ 1-4 0.55 / 0.74 0.54 / 0.41 0.54 / 0.75 0.54 / 0.41 0.55 / 0.75 0.59 / 0.76 0.75 / 0.83
1-1 $\mapsto$ 1-5 0.54 / 0.79 0.55 / 0.84 0.56 / 0.85 0.55 / 0.79 0.65 / 0.87 0.66 / 0.87 0.82 / 0.96
1-1 $\mapsto$ 1-6 0.71 / 0.88 0.71 / 0.91 0.71 / 0.87 0.71 / 0.91 0.44 / 0.84 0.73 / 0.84 0.89 / 0.92
1-1 $\mapsto$ 1-7 0.48 / 0.55 0.48 / 0.50 0.49 / 0.54 0.48 / 0.50 0.31 / 0.57 0.51 / 0.53 0.88 / 0.91
1-1 $\mapsto$ 1-8 0.55 / 0.57 0.53 / 0.70 0.55 / 0.58 0.54 / 0.56 0.52 / 0.56 0.58 / 0.58 0.79 / 0.72
Boiler 1 $\mapsto$ 2 0.44 / 0.64 0.43 / 0.48 0.43 / 0.54 0.43 / 0.48 0.53 / 0.88 0.50 / 0.59 0.61 / 0.66
1 $\mapsto$ 3 0.40 / 0.28 0.40 / 0.11 0.43 / 0.49 0.42 / 0.18 0.41 / 0.48 0.50 / 0.67 0.74 / 0.81
2 $\mapsto$ 1 0.39 / 0.18 0.40 / 0.21 0.40 / 0.36 0.40 / 0.15 0.53 / 0.90 0.51 / 0.66 0.58 / 0.65
2 $\mapsto$ 3 0.40 / 0.38 0.40 / 0.20 0.45 / 0.39 0.42 / 0.21 0.41 / 0.49 0.50 / 0.69 0.59 / 0.63
3 $\mapsto$ 1 0.39 / 0.20 0.40 / 0.16 0.39 / 0.31 0.40 / 0.15 0.48 / 0.67 0.51 / 0.67 0.72 / 0.84
3 $\mapsto$ 2 0.48 / 0.54 0.49 / 0.48 0.49 / 0.55 0.49 / 0.48 0.46 / 0.31 0.50 / 0.57 0.58 / 0.68
Appendix C Extended Experiments
C.1 UDA Comparison
Table 6, adapted from Lai et al. (2023), now includes the results of our model in its final column. This table provides a comprehensive comparison of the performances of various models on the SMD and Boiler datasets, as measured by Macro F1 and AUROC scores.
In this comparative analysis, the DACAD model distinctly outperforms others, consistently achieving the highest scores in both Macro F1 and AUROC across the majority of test cases in both datasets. Its performance is not only superior but also remarkably stable across different scenarios within each dataset, showcasing its robustness and adaptability to diverse data challenges.
Following DACAD, the ContexTDA model frequently ranks as the runner-up, particularly in the SMD dataset, where it often secures the second-highest scores in either or both the Macro F1 and AUROC metrics. Interestingly, certain models exhibit a degree of specialization. For example, SASA, which underperforms in the SMD dataset, demonstrates notably better results in specific scenarios within the Boiler dataset, particularly in terms of the Macro F1 score. This suggests that some models may be more suited to specific types of data or scenarios.
In conclusion, while the DACAD model emerges as the most effective across most scenarios, the varying performances of different models across scenarios and datasets emphasize the need to carefully consider the unique characteristics and capabilities of each model when selecting the most appropriate one for a specific task. This nuanced approach is crucial in leveraging the strengths of each model to achieve optimal results.

### v2

@@ V-C UDA Comparison
To assess our model against ContextDA [13] -the only exiting UDA for the TSAD model-, we use the same datasets and metrics reported in ContextDA’s main paper, as its code is unavailable. On the SMD dataset, ContextDA achieves an average macro F1 of 0.63 and AUROC of 0.75, whereas our model achieves a higher average macro F1 of 0.81 and AUROC of 0.86. Similarly, on the Boiler dataset, ContextDA’s average macro F1 is 0.50 and AUROC 0.65, compared to our model’s superior performance with an average macro F1 of 0.63 and AUROC of 0.71. In this comparative analysis, the DACAD model distinctly outperforms others, consistently achieving the highest scores in both Macro F1 and AUROC across the majority of test cases in both datasets. Its performance is not only superior but also remarkably stable across different scenarios within each dataset, showcasing its robustness and adaptability to diverse data challenges.
The ContexTDA model frequently ranks as the runner-up after DACAD, particularly in the SMD dataset, where it often secures the second-highest scores in either or both the Macro F1 and AUROC metrics. Interestingly, certain models exhibit a degree of specialization. For example, SASA, which underperforms in the SMD dataset, demonstrates notably better results in specific scenarios within the Boiler dataset, particularly in terms of the Macro F1 score. This suggests that some models may be more suited to specific types of data or scenarios.
In conclusion, while the DACAD model emerges as the most effective across most scenarios, the varying performances of different models across scenarios and datasets emphasise the need to carefully consider the unique characteristics and capabilities of each model when selecting the most appropriate one for a specific task. This nuanced approach is crucial in leveraging the strengths of each model to achieve optimal results.
V-D Scalability Test
Figure 2: Comparison of execution time and peak memory usage for DACAD and other models. The top plots show execution time against time series length and dimensionality, while the bottom plots show memory usage.
Scalability tests are crucial to evaluate a model’s performance as data scales up in size and complexity. The plots in Fig.2 provide a comprehensive comparison of DACAD’s scalability with other models described in section V-B. The experiment measures execution time and peak memory usage in relation to the length of time series and their dimensionality. For the dimensionality test, eight synthetic time series datasets with varying dimensions (ranging from 8 to 1024, doubling each step) and a fixed length of 2,000 are generated using the NeurIPS-TS [69]. Another set of eight synthetic datasets with a fixed dimension of 8 varying lengths (ranging from 2,000 to 256,000) are created for the scale-up test concerning length. We report the execution time and memory usage of all models with the same parameter settings as in section V-B.
DACAD demonstrates excellent scalability in terms of time series length, outperforming other domain adaptation models such as VRADA and CLUDA. It maintains lower execution times, particularly evident with larger datasets, highlighting its efficiency. DACAD’s scalability regarding time series dimensions is also remarkable. It remains stable as dimensionality increases, with execution times growing at a very small slope, ensuring it never exceeds reasonable memory limits. Compared to all other anomaly detection models, DACAD shows good scalability concerning time series length, consistently outperforming other methods and demonstrating its robustness in various performance metrics.
V-E Ablation Study
@@
TABLE IV: Effect of loss components on MSL dataset (source: F-5).
@@
Table IV offers several key insights into the impact of different loss components on the MSL dataset using F-5 as a source. Removing the target self-supervised CL (w/o $\mathcal{L}_{\text{SelfCont}}$) leads to lower metrics (F1: 0.481, AUPR: 0.495, AUROC: 0.697). Moreover, excluding source-supervised CL (w/o $\mathcal{L}_{\text{SupCont}}$) reduces effectiveness (F1: 0.463, AUPR: 0.427, AUROC: 0.639), highlighting its role in capturing source-specific features which are crucial for the model’s overall accuracy. Similarly, omitting the discriminator component results in performance reduction (F1: 0.471, AUPR: 0.484, AUROC: 0.699). However, the most significant decline occurs without the classifier (F1: 0.299, AUPR: 0.170, AUROC: 0.503), underscoring its crucial role in effectively distinguishing between normal/anomaly classes. Overall, the best results (F1: 0.595, AUPR: 0.554, AUROC: 0.787) are achieved with all components, highlighting the effectiveness of an integrated approach.
@@
To elucidate the effectiveness of UDA within DACAD, we examine the feature representations from the MSL dataset, as illustrated in Fig. 3. It presents the t-SNE 2D embeddings of DACAD feature representations $\phi^{CL}(\phi^{R}(w))$ for the MSL dataset. Each point represents a time series window, which can be normal, anomalous, or anomaly-injected. These representations highlight the domain discrepancies between source and target entities and demonstrate how DACAD aligns the time series window features effectively. The comparison of feature representations with and without UDA reveals a significant domain shift when UDA is not employed, between source (entity F-5) and the target (entity T-5) within the MSL dataset.
@@
Effect of CEC: Table V compares the performance of our CEC classifier with two BCE and DeepSVDD - on the MSL dataset using F-5 as a source. Our proposed CEC shows superior performance compared to BCE and DeepSVDD across three metrics on the MSL dataset. With the highest F1 score, it demonstrates a better balance of precision and recall. Its leading performance in AUPR indicates greater effectiveness in identifying positive classes in imbalanced datasets. Additionally, CEC’s higher AUROC suggests it is more capable of distinguishing between classes.
TABLE V: Effect of CEC classifier on MSL dataset (source: F-5).
@@
TABLE VI: Effect of anomaly injection on MSL dataset (source: F-5).
@@
Effect of anomaly injection: We study the impact of anomaly injection in Table VI, on the MSL dataset when using F-5 as a source. It shows that anomaly injection significantly improves all metrics (F1: 0.595, AUPR: 0.554, AUROC: 0.787), enhancing the model’s ability to differentiate between normal and anomalous patterns, thereby improving DACAD’s overall accuracy. Without anomaly injection, there is a notable decline in performance, emphasizing its role in precision. The higher standard deviation in AUROC scores without injection suggests more variability and less stability in the model performance. The study underscores the vital role of anomaly injection in improving anomaly detection models. It reveals that incorporating anomaly injection not only boosts detection accuracy but also enhances the model’s overall stability and reliability.
VI Conclusion
@@
The DACAD model is an innovative approach to TSAD that is particularly effective in environments with limited labeled data. By combining domain adaptation and contrastive learning, it applies labelled anomaly data from one domain to detect anomalies in another. Its anomaly injection mechanism, introducing a spectrum of synthetic anomalies, significantly bolsters the model’s adaptability and robustness across various domains. Our evaluations on diverse real-world datasets establish DACAD’s superiority in handling domain shifts and outperforming existing models. Its capability to generalize and accurately detect anomalies, regardless of the scarcity of labeled data in the target domain, represents a significant contribution to TSAD. In future work, we aim to refine the anomaly injection process further, enhancing the model’s ability to simulate a broader range of anomalous patterns. Additionally, we plan to evolve the model to encompass univariate time series analysis, broadening its scope and utility.